Expected manufacturing costs for Crown Point Data Devices are as follows:

Variable Costs Fixed Costs per Month
Direct material $5.00/unit Supervisory salaries $16,000
Direct labor 3.00/unit Factory depreciation 10,700
Variable overhead 1.30/unit Other factory costs 3,250


Estimate manufacturing costs for production levels of 13,000 units, 15,500 units, and 19,000 units per month.

13,000 units 15,500 units 19,000 units
Manufacturing costs $

$

$
Answer #1

Total Variable cost = 5 + 3 + 1.30 = 9.30 per unit

Total fixed cost = 16,000+10,700+3250 = 29,950

13,000 units 15,500 units 19,000 units
Manufacturing cost

= (13,000*9.30) +29950

= 150,850

= (15500*9.30)+29950

= 174,100

= (19,000*9.30)+29950

= 206,650

Comment if you face any issues